# Random-sketching Techniques to Enhance the Numerically Stability of Block Orthogonalization Algorithms for $s$-step GMRES

### v1

@@ 5.2 Randomized-Householder CholQR (RandCholQR)
Figure 4: Randomized Householder CholQR (RandCholQR), where $\mbox{HH}(V)$ returns the orthogonal basis vectors $Q$ and the upper-triangular matrix $R$ based on the Householder QR algorithm such that $V=QR$.
Instead of forming the Gram matrix of the basis vectors, which is the main cause of the numerical instability, RandCholQR first computes the random sketch $\widehat{\mathbf{v}}_{j}$ of the basis vectors $\widehat{V}_{j}$. It then generates their well-conditioned basis vectors $\widehat{Q}_{j}$ using the upper-triangular matrix computed by the stable QR factorization of the sketched vectors $\widehat{\mathbf{v}}_{j}$. When the input vectors $\widehat{V}_{j}$ are numerically full-rank and the sketch matrix $\Theta$ is an $(\varepsilon,\delta,s)$-oblivious $\ell_{2}$-subspace embedding (see Definition 3.2), it can be shown that the generated basis vectors $\widehat{Q}_{j}$ has the condition number of $\mathcal{O}(1)$ [15, 32]. Hence, according to (7), as we call CholQR on $\widehat{Q}_{j}$, the resulting vector $Q_{j}$ has an $\mathcal{O}(\epsilon)$ orthogonality error. Figure 4 shows the pseudocode of the resulting intra-block orthogonalization scheme.
The sketching typically requires one synchronization, and as a result, we expect RandCholQR to perform similarly to CholQR2. The actual performance of the algorithm depends on the type of the sketching being used [12]. In this paper, we look at the following random-sketching techniques that can be implemented using standard linear algebra kernels. Regardless of the types of the sketching used, the last step of the randCholQR requires $\mathcal{O}(ns^{2})$ computation (Line 6 of Figure 4). Hence, in the discussion below, we focus on the first two steps of the algorithm (Lines 2 and 4).
@@
    Gaussian-Sketching can be implemented using a dense GEneral Matrix-Matrix multiply (GEMM). The nice feature of this approach is that it requires the sketch size of only $\mathcal{O}(s)$ [33]. However, this is a “dense” sketch, and the dense sketch matrix $\Theta$ needs to be explicitly stored to call GEMM. Hence, the Gaussia Sketch has the storage overhead of $\mathcal{O}(ns)$ and the computational complexity of $\mathcal{O}(ns^{2})$ to generate the sketch $\widehat{\mathbf{v}}_{j}$, and its overall computational complexity is $\mathcal{O}(ns^{2}+s^{3})$. This overhead could become significant, especially when we need to sketch a large number of basis vectors (e.g., for the two-stage approach discussed in Section 6).
@@
    In addition, the Householder QR factorization of the sketched vectors is performed redundantly on a CPU by each MPI process (Line 4 of Figure 4), leading to a potential parallel performance bottleneck (i.e., the $\mathcal{O}(s^{4})$ complexity for the local computation with the Count-Sketch, compared to the $\mathcal{O}(s^{3})$ complexity with the Gaussian-Sketch).
@@
    where $\Theta_{c}$ and $\Theta_{g}$ are the Count and Gaussian sketch matrices, respectively because the Gaussian sketch is applied locally before the global all-reduce. Hence the communication volume for the all-reduce is $\mathcal{O}(s^{2})$, and is the same as the Gaussian-Sketch and is reduced from $\mathcal{O}(s^{3})$ needed for Count-Sketch.
    Though it is now requires only $\mathcal{O}(s^{3})$ computation to compute the QR factorization of the final sketch, the Gaussian sketch is applied redundantly on each MPI process, which has the computational complexity of $\mathcal{O}(s^{4})$ and is the same as that needed for the Count-Sketch. Nevertheless, the local GEMM for the Count-Gauss may perform more efficiently than the local HH for Count-Sketching, and GEMM before the global-reduce is computed on a GPU, while HH afer the global reduce is computed on CPU.
    Compared to the Gaussian-Sketch, the Count-Gaussian Sketch reduces the computation complexity to apply the sketch from $\mathcal{O}(ns^{2})$ to $\mathcal{O}(ns)$. Moreover, even if the Gaussian-Sketch did not lead to a performance improvement in practice (due to a more efficient dense matrix-matrix multiply implementation, compared to a sparse-matrix matrix multiply), the storage requirement is reduced from $\mathcal{O}(ns)$ to $\mathcal{O}(n)$.
    The overall computational complexity of the Count-Gauss sketch is now $\mathcal{O}(ns+s^{4})$.
@@
When RandCholQR is used as the first intra-block orthogonalization, the orthogonality error of the resulting basis vectors $Q_{1:j}$ is of the order of the machine precision when the following condition is satisfied:
@@
or equivalently if MPK generates numerically full-rank basis vectors $\widehat{V}_{j}$.
The proof of Proposition 5.2 follows by combining the results from two facts: first, BCGS2 attains $O(\epsilon)$ orthogonality error overall provided each “IntraBlk” step in line 3 attains $O(\epsilon)$ orthogonality error [7], and RandCholQR attains $O(\epsilon)$ orthogonality error when (9) is satisfied [32].
Hence, compared to CholQR2, which requires (8), RandCholQR significantly improves the robustness of the overall block orthogonalization process, while we expect only a small overhead in term of its execution time (as shown in Section 11).
@@ 6 Two-stage Block Orthogonalization Framework
BCGS2 discussed in Section 5 can orthogonalize the set of $s+1$ basis vectors with only five synchronizations and using BLAS-3 operations to performs most of its local computation. However, its performance may be still limited by the small step size $s$ required to maintain the stability of MPK. In order to improve the performance of the block orthogonalization while using the small step size, “two-stage” block orthogonalization algorithms have been proposed [6]. Instead of fully-orthogonalizing the basis vectors at every $s$ steps, the two-stage algorithm only “pre-processes” the $s$-step basis vectors $V_{j}$ at every $s$ steps. The objective of this first stage is to maintain the well-conditioning of the basis vectors at a cost that is lower than that required for the full-orthogonalization or with the same cost as the initial orthogonalization (e.g., BCGS with CholQR). Then once a sufficient number of basis vectors, $\widehat{s}$, are generated to obtain higher performance, the second stage orthogonalizes the $\widehat{s}$ basis vectors at once. For the following discussion, we refer to the blocks of $s$ and $\widehat{s}$ vectors as the “panels” and “big panels”, respectively. Also, to distinguish from the two-stage algorithms, we refer to the BCGS algorithms discussed in Section 5 as “one-stage” algorithms.
For this paper, we focus on the specific variant of the two-stage algorithm shown in Figure 5 with MPK. This variant first (roughly) orthogonalizes the new panel basis vectors $V_{j}$ against the previous big panels using BCGS (Line 10). It then pre-processes the resulting panel vectors $\widehat{V}_{j}$ within the current big panel (Line 12). Finally, the second stage fully-orthogonalize the big panel, first orthogonalizing the big panel based on CholQR (Line 15), followed by BCGS with CholQR of the big panels (Line 19).
@@
Without pre-processing step (i.e., one-stage with $s=\widehat{s}$), the condition number of the basis vectors will increase exponentially with the step size (see numerical results in [6]).
As it can be seen in Figure 6, without MPK, this two-stage algorithm can be considered as BCGS2 (shown in Figure 2(c)) that orthogonalizes the big panel as the block vectors at a time, where the pre-processing step, followed by CholQR, is used as the first intra-block orthogonalization. Hence, if the condition number of the big panel $\widehat{Q}_{\ell+1:t+1}$, after the pre-processing step (Lines 4 to 6), is $\mathcal{O}(1)$, then the orthogonality error of $Q_{\ell+1:t+1}$ after the first CholQR (Line 7) is $\mathcal{O}(\epsilon)$, and thus, the overall two-stage block orthogonalization is stable with the $\mathcal{O}(\epsilon)$ orthogonality error of the resulting vectors $Q_{1:t+1}$. In the next section, we introduce two pre-processing schemes and discuss the condition on the input big panel $\kappa(\widehat{V}_{\ell+1:t+1})$ to ensure the overall stability of the two-stage algorithm.
@@ 7.1 BCGS with Pythagorean Inner Product (BCGS-PIP)
Our first pre-processing scheme for the two-stage algorithm is based on the single-reduce variant [29, 35] of BCGS with CholQR, shown in Figure 7. Instead of explicitly computing the Gram matrix, this variant uses the Pythagorean rule, hence requiring only one global-reduce for both the inter-block and intra-block orthogonalization.111 For the first block (i.e., $j=\ell+1$), BCGS-PIP is equivalent to CholQR. In addition, it was shown [29, Theorem 3.4] that if the condition number of the input basis vectors is bounded as
@@
where $t+1$ is the last block index of the big panel (i.e., $t:=\ell+\widehat{s}/s-1$), then the orthogonality error of the big panel computed by BCGS-PIP satisfies
$\|I-\widehat{Q}_{\ell+1:t+1}^{T}\widehat{Q}_{\ell+1:t+1}\|\leq c_{4}(\epsilon)\kappa(\widehat{V}_{\ell+1:t+1})^{2}.$ (12)
@@
By [7], it is sufficient to prove that the big panel $Q_{\ell+1:t+1}$ before the reorthogonalization (after Line 7 of Figure 6) satisfies
$\|I-Q_{\ell+1:t+1}^{T}Q_{\ell+1:t+1}\|=O(\epsilon).$
To prove this, because of the assumption (11) and the bound (12), we note that
$\|I-\widehat{Q}^{T}_{\ell+1:t+1}\widehat{Q}_{\ell+1:t+1}\|=O(\epsilon)\kappa^{2}\left(\widehat{V}_{\ell+1:t+1}\right)=O(1),$
and thus by Weyl’s inequality, we have
$\kappa\left(\widehat{Q}_{\ell+1:t+1}\right)=O(1).$ (13)
Hence, with the bound (7), the condition (13) implies that $Q_{\ell+1:t+1}$ produced by CholQR of $\widehat{Q}_{\ell+1:t+1}$ at Line 7 satisfies
$\|I-Q_{\ell+1:t+1}^{T}Q_{\ell+1:t+1}\|=O(\epsilon),$
and the proof follows.
@@
Similarly to CholQR, BCGS-PIP can fail when the condition number of the big panel is greater than the reciprocal of the square-root of the machine epsilon $\epsilon$. This can cause numerical issues.
@@ 7.2 Randomized BCGS
1:Input: $V_{j}$ (new block), 2:           $Q_{j_{1}:j-1}$ (previously-orthogonalized blocks), and 3:           $\mathbf{q}_{j_{1}:j-1}$ (previous sketch vectors) 4:Output: $Q_{j}$, $\mathbf{q}_{j}$, $R_{1:j,j}$ 5:// Sketch the new “panel” 6:$\mathbf{v}_{j}:=\Theta^{T}V_{j}$ 7:// Orthogonalize new sketched panel within big panel 8:[$\mathbf{q}_{j}$, $R_{j_{1}:j,j}$] := BCGS2-HH($\mathbf{q}_{j_{1}:j-1}$, $\mathbf{v}_{j}$) 9:// Generate well-conditioned basis 10:$\widehat{V}_{j}:=V_{j}-Q_{j_{1}:j-1}R_{j_{1}:j-1,j}$ 11:$Q_{j}:=\widehat{V}_{j}R^{-1}_{j,j}$
Figure 8: Randomized BCGS (RandBCGS).
To enhance the stability of the pre-processing scheme based on BCGS-PIP (especially for the ill-conditioned basis vectors generated by MPK), we consider a randomized BCGS scheme shown in Figure 8 as our pre-processing algorithm. This algorithm sketches the big panel, but $s$ basis vectors at a time. In order to ensure the stability, we orthogonalize the sketched vectors $\widehat{\mathbf{v}}_{\ell+1:t+1}$ using BCGS2 with Householder intra-block orthogonalization. Similar randomized BCGS algorithms were discussed in [14]. We used this randomized algorithm as the pre-processing scheme for the two-satage BCGS2, in order to maintain the well-conditioning of the big panel $\widehat{Q}_{\ell+1:t+}$ and obtain the overall $\mathcal{O}(\epsilon)$ orthogonality error of $Q_{1:t+1}$.
Conjecture 7.1.
The overall stability of two-stage algorithm with randomized BCGS pre-processing is ensured when
@@
or equivalently if MPK generates numerically full-rank basis vectors for each big panel $\widehat{V}_{\ell+1:t+1}$.
Assuming the condition (10) is satisfied, the condition (14) is equivalent to the condition (9), and hence the two-stage algorithm with the randomized BCGS is as stable as the one-stage algorithm with random sketching.
@@
BCGS-PIP and RandBCGS, shown in Figures 7 and  8, respectively, have a similar algorithmic structure. In particular, both algorithms would require one global-reduce, followed by a small local computation (either Cholesky factorization of the Gram matrix or BCGS2 orthogonalization of the sketched vectors) and backward-substitution to generate the basis vectors $\widehat{Q}_{j}$. Though RandBCGS has the larger complexity for the local computation, the main factor that impacts their performance difference is the dot-products required for BCGS-PIP and the random-sketching required for RandGCGS.

### v2

@@ 5.2 Randomized-Householder CholQR (RandCholQR)
1:Input: $V_{j}$ 2:Output: $\widehat{V}_{j}$, $R_{j,j}$ 3:// Sketch the new “panel” 4:$\mathbf{v_{j}}:=\Theta^{T}V_{j}$ 5:// Orthogonalize new sketched panel 6:$[\mathbf{q}_{j},R_{j,j}]:=\mbox{HH}(\mathbf{v}_{j})$ 7:// Generate well-conditioned basis 8:$\widehat{V}_{j}:=V_{j}R_{j,j}^{-1}$
(a) Randomized Householder QR (RandHH)
@@
(b) Randomized Householder CholQR (RandCholQR)
Figure 4: Randomized QR algorithm, where $\mbox{HH}(V)$ returns the orthogonal basis vectors $Q$ and the upper-triangular matrix $R$ based on the Householder QR algorithm such that $V=QR$.
Instead of forming the Gram matrix of the basis vectors, which is the main cause of the numerical instability, Randomized CholQR (RandCholQR) first computes the random sketch $\widehat{\mathbf{v}}_{j}$ of the basis vectors $\widehat{V}_{j}$. It then generates their well-conditioned basis vectors $\widehat{Q}_{j}$ using the upper-triangular matrix computed by the stable QR factorization of the sketched vectors $\widehat{\mathbf{v}}_{j}$. When the input vectors $\widehat{V}_{j}$ are numerically full-rank and the sketch matrix $\Theta$ is a $(\mu,\delta,s)$-oblivious $\ell_{2}$-subspace embedding (see Definition 3.2), it can be shown that the generated basis vectors $\widehat{Q}_{j}$ has the condition number of $\mathcal{O}(1)$ [14, 33]. Hence, according to (7), as we call CholQR on $\widehat{Q}_{j}$, the resulting vector $Q_{j}$ has an $\mathcal{O}(\epsilon)$ orthogonality error. Figure 4(b) shows the pseudocode of the resulting RandCholQR algorithm for the intra-block orthogonalization.
The sketching typically requires one synchronization, and as a result, we expect RandCholQR to perform similarly to CholQR2. The actual performance of the algorithm depends on the type of the sketching being used. In this paper, we look at the following random-sketching techniques that can be implemented using standard linear algebra kernels. Regardless of the types of the sketching used, the last step of the randCholQR, to generate the well-conditioned basis vectors through forward substitutions, requires $\mathcal{O}(ns^{2})$ computation (Line 6 of Figure 4(b)). Hence, in the discussion below, we focus on the first two steps of the algorithm (Lines 2 and 4).
@@
    Gaussian-Sketching can be implemented using a dense GEneral Matrix-Matrix multiply (GEMM). The nice feature of this approach is that it requires the sketch size of only $\mathcal{O}(s)$ [23]. However, this is a “dense” sketch, and the dense sketch matrix $\Theta$ needs to be explicitly stored to call GEMM. Hence, the Gaussia Sketch has the storage overhead of $\mathcal{O}(ns)$ and the computational complexity of $\mathcal{O}(ns^{2})$ to generate the sketch $\widehat{\mathbf{v}}_{j}$, and its overall computational complexity is $\mathcal{O}(ns^{2}+s^{3})$.
    Though this complexity is the same as CholQR, the complexity of RandHH with the dense Gaussian-Sketch has a larger constant associated with the sketch size (i.e., $ns^{2}$ flops for CholQR to compute the Gram matrix, compared to $4ns^{2}$ floating-point operations (flops) for RandHH to generate the sketched vectors with the sketch size of $2s$).
    The terms associated with $s^{2}$ and $s^{3}$ in the complexity $\mathcal{O}(ns^{2}+s^{3})$ could become significant, especially when we need to sketch a large number of basis vectors (e.g., for the two-stage approach discussed in Section 6).
@@
    In addition, the Householder QR factorization of the sketched vectors is performed redundantly on a CPU by each MPI process (Line 4 of Figure 4(b)), leading to a potential parallel performance bottleneck (i.e., the $\mathcal{O}(s^{4})$ complexity for the local computation with the Count-Sketch, compared to the $\mathcal{O}(s^{3})$ complexity with the Gaussian-Sketch).
@@
    where $\Theta_{c}$ and $\Theta_{g}$ are the Count and Gaussian sketch matrices, respectively. The communication volume is reduced from the Count-Sketch because the Gaussian sketch is applied locally before the global all-reduce. Hence the communication volume for the all-reduce is $\mathcal{O}(s^{2})$, and is the same as the Gaussian-Sketch and is reduced from $\mathcal{O}(s^{3})$ needed for Count-Sketch.
    Though it is now requires only $\mathcal{O}(s^{3})$ computation to compute the QR factorization of the final sketch, the Gaussian sketch is applied redundantly on each MPI process, which has the computational complexity of $\mathcal{O}(s^{4})$ and is the same as that needed for the Count-Sketch. Nevertheless, the local GEMM for the Count-Gauss may perform more efficiently than the local HH for Count-Sketching. Moreover, in our implementation, GEMM before the global-reduce is computed on a GPU, while HH afer the global reduce is computed on CPU.
    Compared to the Gaussian-Sketch (or to the generation of the Gram matrix for CholQR), the Count-Gaussian reduces the computation complexity to apply the sketch from $\mathcal{O}(ns^{2})$ to $\mathcal{O}(ns)$. Moreover, even if the Count-Gaussian did not lead to a performance improvement over the Gaussian-Sketch in practice (due to a more efficient dense matrix-matrix multiply implementation, compared to a sparse-matrix matrix multiply), the storage requirement is reduced from $\mathcal{O}(ns)$ to $\mathcal{O}(n)$.
    The overall computational complexity of the Count-Gauss is $\mathcal{O}(ns+s^{4})$.
@@
When RandCholQR is used as the first intra-block orthogonalization, the orthogonality error of the resulting basis vectors $Q_{1:j}$ is of the order of the machine precision when $\Theta^{T}$ forms an $\mu$-subspace embedding over the range of $\widehat{V}_{j}$, and the following condition is satisfied:
@@
where $c_{2}(\epsilon,n,s)$ is a constant, or equivalently if MPK generates numerically full-rank basis vectors $\widehat{V}_{j}$.
Proof.
The proof of Proposition 5.2 follows by combining the results from two facts: first, BCGS2 attains $O(\epsilon)$ orthogonality error overall provided each “IntraBlk” step in line 3 attains $O(\epsilon)$ orthogonality error [7], and RandCholQR attains $O(\epsilon)$ orthogonality error when (9) is satisfied [33]. ∎
Hence, compared to CholQR2, which requires (8), RandCholQR improves the numerical robustness of the overall block orthogonalization process, while we expect only a small overhead in term of its execution time (as shown in Section 11).
@@ 6 Two-stage Block Orthogonalization Framework
BCGS2 discussed in Section 5 can orthogonalize the set of $s+1$ basis vectors with only five synchronizations and using BLAS-3 operations to performs most of its local computation. However, its performance may be still limited by the small step size $s$ required to maintain the stability of MPK. In order to improve the performance of the block orthogonalization while using the small step size, “two-stage” block orthogonalization algorithms have been proposed [6]. Instead of fully-orthogonalizing the basis vectors at every $s$ steps, the two-stage algorithm only “pre-processes” the $s$-step basis vectors $V_{j}$ at every $s$ steps. The objective of this first stage is to maintain the well-conditioning of the basis vectors at a cost that is lower than that required for the full orthogonalization or with the same cost as the initial orthogonalization (e.g., BCGS with CholQR). Then once a sufficient number of basis vectors, $\widehat{s}$, are generated to obtain higher performance, the second stage orthogonalizes the $\widehat{s}$ basis vectors at once. For the following discussion, we refer to the blocks of $s$ and $\widehat{s}$ vectors as the “panels” and “big panels”, respectively. Also, to distinguish from the two-stage algorithms, we refer to the BCGS algorithms discussed in Section 5 as “one-stage” algorithms.
For this paper, we focus on the specific variant of the two-stage algorithm shown in Figure 5. This variant first (roughly) orthogonalizes the new panel of basis vectors $V_{j}$ against the previous big panels using BCGS (Line 10). It then pre-processes the resulting panel vectors $\widehat{V}_{j}$ within the current big panel (Line 12). Finally, the second stage fully orthogonalizes the big panel, first orthogonalizing the big panel using CholQR (Line 15), followed by BCGS with CholQR on the big panels (Line 19).
@@
Without the pre-processing step (i.e., one-stage with $s=\widehat{s}$), the condition number of the basis vectors will increase exponentially with the step size (see numerical results in [6]).
As it can be seen as the generic block orthogonalization scheme, without MPK, in Figure 6, this two-stage algorithm can be considered as BCGS2 (shown in Figure 2(c)) that orthogonalizes the big panel as a set of block vectors, where the pre-processing step, followed by CholQR, is used as the first intra-block orthogonalization. Hence, if the condition number of the big panel $\widehat{Q}_{\ell+1:t+1}$, after the pre-processing step (Lines 4 to 6) is $\mathcal{O}(1)$, then the orthogonality error of the big panel $Q_{\ell+1:t+1}$ after the first CholQR (Line 7) is $\mathcal{O}(\epsilon)$, and thus, the overall two-stage block orthogonalization is stable with $\mathcal{O}(\epsilon)$ orthogonality error of the resulting vectors, i.e.,
$\|I-Q_{1:t+1}^{T}Q_{1:t+1}\|=\mathcal{O}(\epsilon).$
In the next section, we introduce two pre-processing schemes and discuss the condition on the input big panel $\kappa(\widehat{V}_{\ell+1:t+1})$ to ensure the overall stability of the two-stage algorithm.
@@ 7 Preprocessing Schemes for Two-stage Framework
We first formally establish the following proposition.
Proposition 7.1.
If a pre-processing scheme can maintain the condition number of the big panel to be bounded as
$\kappa\left(\widehat{Q}_{\ell+1:t+1}\right)=O(1),$ (11)
then the overall stability of the two-stage scheme is ensured.
Proof.
Using (7), condition (11) implies that $Q_{\ell+1:t+1}$ produced by CholQR of $\widehat{Q}_{\ell+1:t+1}$ at Line 7 of Figure 6 satisfies
$\|I-Q_{\ell+1:t+1}^{T}Q_{\ell+1:t+1}\|=O(\epsilon).$
Therefore, by [7], the overall stability of the two-stage scheme is ensured. ∎
@@ 7.1 BCGS with Pythagorean Inner Product (BCGS-PIP)
Our first pre-processing scheme for the two-stage algorithm is based on the single-reduce variant [30, 35] of BCGS with CholQR, shown in Figure 7. Instead of explicitly computing the Gram matrix, this variant uses the Pythagorean rule, hence requiring only one global-reduce for both the inter-block and intra-block orthogonalization.111 For the first block (i.e., $j=\ell+1$), BCGS-PIP is equivalent to CholQR. In addition, it was shown [30, Theorem 3.4] that if the condition number of the input basis vectors is bounded as
@@
where $c_{3}(\epsilon,n,s)$ is a constant and $t+1$ is the last block index of the big panel (i.e., $t:=\ell+\widehat{s}/s-1$), then the orthogonality error of the big panel computed by BCGS-PIP satisfies
$\|I-\widehat{Q}_{\ell+1:t+1}^{T}\widehat{Q}_{\ell+1:t+1}\|\leq c_{3}(\epsilon,n,\widehat{s})\kappa(\widehat{V}_{\ell+1:t+1})^{2}.$ (13)
@@
Because of the assumption (12) and the bound (13), we have
$\|I-\widehat{Q}^{T}_{\ell+1:t+1}\widehat{Q}_{\ell+1:t+1}\|<1/2.$
Thus, by Weyl’s inequality, we obtain
$\left\{\begin{array}[]{l}\sigma_{\min}\left(\widehat{Q}^{T}_{\ell+1:t+1}\widehat{Q}_{\ell+1:t+1}\right)\geq 1-\|I-\widehat{Q}^{T}_{\ell+1:t+1}\widehat{Q}_{\ell+1:t+1}\|>1/2\\ \sigma_{\max}\left(\widehat{Q}^{T}_{\ell+1:t+1}\widehat{Q}_{\ell+1:t+1}\right)\leq 1+\|I-\widehat{Q}^{T}_{\ell+1:t+1}\widehat{Q}_{\ell+1:t+1}\|<3/2.\end{array}\right.$
Therefore,
$\kappa\left(\widehat{Q}_{\ell+1:t+1}\right)=\sqrt{\kappa\left(\widehat{Q}_{\ell+1:t+1}^{T}\widehat{Q}_{\ell+1:t+1}\right)}=O(1),$ (14)
and the proof follows by Proposition 7.1. ∎
@@
Similarly to CholQR, BCGS-PIP can fail when the condition number of the big panel is greater than the reciprocal of the square-root of the machine epsilon $\epsilon$. This can cause numerical issues (especially for the ill-conditioned basis vectors generated by MPK).
@@ 7.2 Randomized BCGS
1:Input: $V_{j}$ (new block), 2:           $\widehat{Q}_{j_{1}:j-1}$ (previous approximately orthogonal blocks), and 3:           $\mathbf{q}_{j_{1}:j-1}$ (previous sketch vectors) 4:Output: $\widehat{Q}_{j}$, $\mathbf{q}_{j}$, $R_{1:j,j}$ 5:// Sketch the new “panel” 6:$\mathbf{v}_{j}:=\Theta^{T}V_{j}$ 7:// Orthogonalize new sketched panel within big panel 8:[$\mathbf{q}_{j}$, $R_{j_{1}:j,j}$] := BCGS2-HH($\mathbf{q}_{j_{1}:j-1}$, $\mathbf{v}_{j}$) 9:// Generate well-conditioned basis 10:$\widehat{V}_{j}:=V_{j}-\widehat{Q}_{j_{1}:j-1}R_{j_{1}:j-1,j}$ 11:$\widehat{Q}_{j}:=\widehat{V}_{j}R^{-1}_{j,j}$
(a) The $j$th step (for PreProc for Figure 6)
1:Input: $V_{1:t+1}$ 2:Output: $\widehat{Q}_{1:t+1}$, $R_{1:t+1,1:t+1}$ 3:// Sketch the big panel 4:$\mathbf{v}_{1:t+1}:=\Theta^{T}V_{1:t+1}$ 5:// Orthogonalize the sketch of the big panel 6:for $j=1,2,\dots,t+1$ do 7:    [$\mathbf{q}_{j}$, $R_{1:j,j}$] := BCGS2-HH($\mathbf{q}_{1:j-1}$, $\mathbf{v}_{j}$) 8:end for 9:// Generate well-conditioned basis by forward-substitution 10:for $j=1,2,\dots,t+1$ do 11:    $\widehat{V}_{j}:=V_{j}-\widehat{Q}_{1:j-1}R_{1:j-1,j}$ 12:    $\widehat{Q}_{j}:=\widehat{V}_{j}R^{-1}_{j,j}$ 13:end for
(b) Accumulated steps.
Figure 8: Randomized BCGS2 (RandBCGS2).
To enhance the stability of the pre-processing scheme based on BCGS-PIP, we consider a randomized BCGS scheme shown in Figure 8 as our pre-processing algorithm. This algorithm sketches the big panel, but $s$ basis vectors at a time. To ensure stability, we orthogonalize the sketched vectors $\widehat{\mathbf{v}}_{\ell+1:t+1}$ using BCGS2 with Householder intra-block orthogonalization. Similar randomized block orthogonalization algorithms were discussed in [14, 15, 33]. We used this randomized algorithm as the pre-processing scheme for the two-satage BCGS2, in order to maintain the well-conditioning of the big panel $\widehat{Q}_{\ell+1:t+1}$ and obtain the overall $\mathcal{O}(\epsilon)$ orthogonality error of $Q_{1:t+1}$.
Proposition 7.3.
The overall stability of two-stage algorithm is ensured when randomized BCGS pre-processing is used with $\Theta^{T}$ that forms an $\mu$-subspace embedding over the range of $\widehat{V}_{\ell+1:t+1}$, and the following condition is satisfied:
@@
where $c_{4}(\epsilon,n,s)$ is a constant, or equivalently when MPK generates numerically full-rank basis vectors for each big panel $\widehat{V}_{\ell+1:t+1}$.
Proof.
Observe that RandBCGS2 (in Figure 8(b)) is identical to RandHH (in Figure 4(a)) except that BCGS2-HH is used, instead of HH factorization, on the sketched vectors to generate the upper-triangular matrix. In [33, Corollary 5.1], it was proven that RandHH results in $\kappa\left(\widehat{Q}_{\ell+1:t+1}\right)=O(1)$.
We prove in Appendix that the backward error of the QR factorization via BCGS2-HH only differs from the Householder QR backward error by a constant factor. Hence, the backward error analysis in [33, Section 5.2.4] only differs by a constant factor when BCGS2+HH, instead of HH, was used on the sketched vectors. Therefore, by [33, Corollary 5.1], RandBCGS2 will generate the basis vectors $\widehat{Q}_{\ell+1:t+1}$ whose condition number is bounded by $\kappa\left(\widehat{Q}_{\ell+1:t+1}\right)=O(1)$, and the proof follows by Proposition 7.1. ∎
Assuming that the condition number of the big panel is in the same order as that of each panel in the one-stage algorithm (as in the condition (10)), the condition (15) on the big panel for the two-stage algorithm is equivalent to the condition (9) on the panel for the one-stage algorithm, and hence the two-stage algorithm with RandBCGS2 is as stable as the one-stage algorithm with RandHH.
The roundoff error analysis of a randomized BCGS, that is similar to the one in Figure 8, has been presented in [14], where their framework allows generating the sketches of the block for the inter and intra block orthogonalization, separately, while in this paper, we focus on the one in Figure 8 that uses a single sketch of the big panel (similar to randCholQR, but by generating the sketch of each block at a time).
@@
BCGS-PIP and RandBCGS2, shown in Figures 7 and  8, respectively, have a similar algorithmic structure. In particular, both algorithms would require one global-reduce, followed by a small local computation (either Cholesky factorization of the Gram matrix or BCGS2 orthogonalization of the sketched vectors) and forward-substitution to generate the basis vectors $\widehat{Q}_{j}$. Though RandBCGS2 has the larger complexity for the local computation, the main factor that impacts their performance difference is the dot-products required for BCGS-PIP and the random-sketching required for RandGCGS.